### v1

# DOCE: Finding the Sweet Spot for Execution-Based Code Generation

Abstract
Recently, a diverse set of decoding and reranking procedures have been shown effective for LLM-based code generation. However, a comprehensive framework that links and experimentally compares these methods is missing. We address this by proposing Decoding Objectives for Code Execution, a comprehensive framework that includes candidate generation, $n$-best reranking, minimum Bayes risk (MBR) decoding, and self-debugging as the core components. We then study the contributions of these components through execution-based evaluation metrics. Our findings highlight the importance of execution-based methods and the difference gap between execution-based and execution-free methods. Furthermore, we assess the impact of filtering based on trial unit tests, a simple and effective strategy that has been often overlooked in prior works. We also propose self-debugging on multiple candidates, obtaining state-of-the-art performance on reranking for code generation. We expect our framework to provide a solid guideline for future research on code generation. [^1] Our code is available at https://github.com/deep-spin/doce.
DOCE: Finding the Sweet Spot for Execution-Based Code Generation
Haau-Sing Li1,2, Patrick Fernandes2,3,4, Iryna Gurevych1, André F. T. Martins2,4,5, 1UKP Lab, Department of Computer Science and Hessian Center for AI (hessian.AI), TU Darmstadt 2Instituto de Telecomunicações 3Language Technologies Institute, Carnegie Mellon University 4Instituto Superior Técnico, Universidade de Lisboa & ELLIS Unit Lisbon 5Unbabel Correspondence: hli@ukp.tu-darmstadt.de
1 Introduction
Despite the impressive capabilities of large language models (LLMs) in code generation (Chen et al., 2021; Li et al., 2022, 2023; Rozière et al., 2024; Luo et al., 2024; Guo et al., 2024a; Lozhkov et al., 2024), demonstrated by their strong performance on popular execution-based metrics (Hendrycks et al., 2021; Chen et al., 2021; Austin et al., 2021; Lai et al., 2022), the predominant methods for generating code answers still rely on deterministic decoding techniques, such as greedy decoding and beam search, which can lead to suboptimal output (Shi et al., 2022).
Figure 1: Under the best settings, our proposed framework shows considerable improvement in execution accuracy over greedy decoding. We show results using CodeLlama-{7,13}B-Instruct (Rozière et al., 2024) and DeepSeekCoder-6.7B-Instruct (Guo et al., 2024a).
An alternative approach involves generating multiple candidates through sampling a generative model and selecting one using $n$-best reranking (Ng et al., 2019; Bhattacharyya et al., 2021; Cobbe et al., 2021; Shen et al., 2021) or Minimum Bayes Risk (MBR) decoding (Eikema and Aziz, 2020, 2022). Metrics for reranking have been extensively explored in domains like machine translation (Eikema and Aziz, 2020; Freitag et al., 2022; Farinhas et al., 2023), where it has been shown to outperform greedy decoding and beam search. Similar methods in code generation include reranking with likelihoods of generated programs and instructions (Zhang et al., 2022) and MBR decoding with execution outputs (Shi et al., 2022). However, a unified framework providing an apples-to-apples comparison between these methods is currently missing.
Figure 2: The Decoding Objectives for Code Execution (DOCE) Framework. Firstly, multiple candidates are generated through sampling. Each candidate then is assigned a score using an $n$-best reranker or MBR, before the candidate with the highest score is returned. Optionally, self-debugging can be applied to multiple candidates before scoring as we propose, or the highest score candidate as proposed by Chen et al. (2024).
Bridging this gap is especially significant for current research. First of all, with lately released open-source LLMs like CodeLlama (Rozière et al., 2024), there is demand for understanding the effects of reranking methods. For example, filtering based on trial unit tests (Li et al., 2022) was initially proposed when an LLM can generate few sensible code candidates, while lately released LLMs like CodeLlama (Rozière et al., 2024) can generate sensible programs easily.Additionally, evaluation with substantially more high-quality unit tests (Liu et al., 2023) is more rigorous and can potentially boost execution-based MBR decoding (Li et al., 2022; Shi et al., 2022). Last but not least, more metrics continue to emerge (Zhang et al., 2022; Zhou et al., 2023; Dong et al., 2023), providing an opportunity to comprehensively understand the execution component of reranking and self-debugging compared with execution-free approaches.
Therefore, we propose a unified framework for execution-based code generation to improve decoding performance on execution-based metrics (see Figure 1). Our contributions are listed below:
[…]
  • We analyze the impact of the number of generated candidates and sampling temperature. Specifically, sampling with previously unseen high temperatures allows high oracle reranking performance and execution-based MBR performance with filtering based on trial unit tests.
  • With analysis of different approaches, we emphasize the importance of execution-based approaches in the context of candidate generation, reranking, and self-debugging. Specifically, our analysis reveals the importance of filtering based on trial unit tests, a commonly used technique whose effect has not been reported in previous works.
  • By applying DOCE to state-of-the-art language models, we find that execution-based reranking yields close-to-oracle performance. We show that self-debugging, a technique applied by Chen et al. (2024) as a final step, is more effective and robust when applied to multiple generated candidates before the final choice is made.
2 Candidate Generation, Self-Debugging, and Reranking
2.1 Candidate generation
An LLM-based code generation model defines a probability distribution $p_{\theta_{\mathrm{LLM}}}(y|x)$ over the space of possible programs, conditioned on a description of the task $x$, where $\theta_{\mathrm{LLM}}$ refers to learned parameters of the LLM. An $x$ often contains a natural language description; some unit test cases, namely trial unit tests, are used to regularize and generate the corresponding code, and the function heads optionally.
Generated code is usually predicted using greedy decoding or $p$-nucleus sampling (Holtzman et al., 2020). While greedy decoding generates a single high-probability candidate, sampling allows obtaining a set of diverse generated candidates $\mathcal{Y}$. Also note that in practice, nucleus sampling is often accompanied by temperature sampling (Dabre and Fujita, 2021). Previous studies (Shi et al., 2022; Chen et al., 2023) also found that sampling-based methods yield higher oracle performance (the oracle on models’ performance, choosing the optimal candidate) denoted by pass@k where k is the number of candidates generated). However, it is still unclear how to best pick a single candidate from the set of candidates.
2.2 Reranking
We now assume access to $\mathcal{Y}$ containing $N$ generated candidates for a given $x$, as mentioned in Section 2.1. To provide a final output $\hat{y}$, we explore how to rank and pick candidates for code generation effectively. Note that with outputs that have the same score, we choose the output with the smallest index.
2.2.1 $n$-Best Reranking.
In its simplest form, $n$-best reranking uses a single feature $f$ and ranks candidates according to it, with the chosen candidate then being the one that maximizes this feature.
$\hat{y}=\operatorname*{argmax}_{y^{\prime}\in\mathcal{Y}}f(y^{\prime}).$ (1)
Execution-based Feature.
[…]
$f(y^{\prime})=\prod_{t\in T_{\text{trial}}}\mathds{1}_{t(y^{\prime})=\text{pass}}.$ (2)
Scores from External Models.
Other than execution-based features, we can use a single model to predict the quality score of generated code with a trained model $\theta$. Specifically, previous efforts include training models using execution success/failure as signals (Inala et al., 2022; Ni et al., 2023), given the input $x$ and generated candidate $y^{\prime}$. During inference, a score $f_{\theta}(x,y^{\prime})$ is given for every generated candidate. Notice that for fair comparison of execution-based and execution-less methods, we do not consider models that require information from execution during inference time (Ni et al., 2023).
2.2.2 MBR Decoding
While $n$-best list reranking can rank candidates based on reference-free features, it is also possible to rank candidates based on reference-based metrics using MBR decoding. Concretely, if we consider a utility function $U(y^{\star},y)$ that measures the similarity between a candidate $y$ and some reference (correct) code $y^{\star}$, MBR selects the candidate in $\mathcal{Y}$ that maximizes the expected utility, when considering all candidates as possible references:
$\hat{y}=\operatorname*{argmax}_{y^{\prime}\in\mathcal{Y}}\frac{1}{|\mathcal{Y}|}\sum_{y\in\mathcal{Y}}U(y,y^{\prime}).$ (3)
While MBR is most frequently used to optimize translation quality metrics like BLEU (Papineni et al., 2002) or COMET (Rei et al., 2020), we are interested in metrics that specifically capture the quality of code. Shi et al. (2022) showed that translation-quality metrics do not outperform greedy decoding when evaluated with execution, for which we do not consider using translation-quality metrics like BLEU.
Execution-based Metrics.
If we can feed unit test inputs to candidates, a basic utility function (Shi et al., 2022; Li et al., 2022) is to consider if a candidate matches the outputs of some reference code with the set of unit tests used for evaluation $\mathcal{E}_{\mathcal{T}}$, the utility function is the exact match of outputs between the reference and the candidates
$U(y,y^{\prime})=\mathds{1}_{\mathcal{E}_{\mathcal{T}}(y)=\mathcal{E}_{\mathcal{T}}(y^{\prime})}.$ (4)
In practice, this utility function reduces the MBR decision to a majority vote for the candidate that produces outputs that agree the most with other candidates’ outputs (Wang et al., 2023).
Note that in Shi et al. (2022) execution-based MBR both on the cases of using only inputs from trial unit tests only and from all unit tests used for evaluation. Since we decide to use these unit tests as filtering, we use all unit tests given but only their inputs. Our practice aligns with most previous common practices (Shi et al., 2022; Ni et al., 2023; Chen et al., 2024).
External Models for MBR.
We can also use external models for MBR. Existing external models include CodeBertScore (Zhou et al., 2023), CodeScore (Dong et al., 2023) to calculate $U(y,y^{\prime})$. Specifically, with a trained model $\theta$, we calculate $U_{\theta}(x,y,y^{\prime})$ by inputting input $x$ and two generated candidates $y$ and $y^{\prime}$.
2.3 Self-Debugging
The performance of reranking is oracleed by the maximum quality of generated candidates. An approach to further improve the quality of generated code is to debug it using the same model used for generation (Chen et al., 2024). Given input $x$, generation $y^{\prime}$, and optionally information of execution on trial unit tests $T_{\text{trial}}$, we obtain a new generation $y^{\prime}_{d}$ using the code LLM parameterized by $\theta_{\mathrm{LLM}}$:
$y^{\prime}_{d}=\mathrm{debug}_{\theta_{\mathrm{LLM}}}(x,y^{\prime},T_{\text{trial}}),$ (5)
where $y^{\prime}_{d}$ represents a corrected version $y$. In practice, this self-debugging can be achieved by prompting the model to correct errors in the presented code, potentially based on the results from execution. We use greedy decoding to get the debugged code, same as Chen et al. (2024).
2.4 DOCE Framework
We define DOCE framework. Firstly, with only reranking and MBR, we have
$\hat{y}=\operatorname*{argmax}_{y^{\prime}\in\mathcal{Y}}\frac{f(y^{\prime})}{|\mathcal{Y}|}\sum_{y\in\mathcal{Y}}U(y,y^{\prime}).$ (6)
Self-Debugging with Single Selected Candidate from Reranking.
If we only debug selected candidates from reranking, namely SD-1 (Chen et al., 2024), we have the final code answer $\textit{debug}_{\theta}(x,\hat{y},T_{\text{trial}})$.
Self-Debugging on All Candidates before Reranking.
We can also perform self-debugging on all generated candidates before reranking, namely SD-Multi. We presume this outperforms self-debugging on one candidate since self-debugging aims to regenerate candidates that pass the trial unit tests, and more candidates passing trial unit tests leads to more candidates for effective MBR decoding with possibly better reranking results. Therefore, the final code answer is $\operatorname*{argmax}_{y^{\prime}_{d}\in\mathcal{Y}_{d}}\frac{f(y^{\prime}_{d})}{|\mathcal{Y}_{d}|}\sum_{y_{d}\in\mathcal{Y}_{d}}U(y_{d},y^{\prime}_{d})$.
3 Experiments
3.1 Datasets
We conduct experiments using two widely recognized execution-based datasets: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). We employ the sanitized, higher-quality subset of the MBPP dataset, namely MBPP-S. Note that HumanEval is a test-only dataset and MBPP-S are normally treated as test-only. We utilize EvalPlus (Liu et al., 2023) which includes more than 35 times unit tests than the original benchmark for both HumanEval and MBPP-S, ensuring a more rigorous assessment of function correctness.
3.2 Candidate Generation
We generate candidates with the CodeLlama-{7B,13B}-Instruct (Rozière et al., 2024) models and DeepSeekCoder-6.7B-Instruct (Guo et al., 2024a) model because instruction fine-tuning allows them to generate output in the markdown format with code blocks bracketed by ```python and ```, allowing self-debugged code to be processed easily. We generate {1,5,10,20,25,50} [^3] Note that we use candidate obtained with greedy decoding when with one generated candidate. candidates for each task. We first explore sampling temperature by varying it between 0.2 and 2.0 with $p$-nucleus sampling $p$ 0.95, before deciding what sampling temperature we use for subsequent experiments. With MBPP-S, we vary temperature between 0.2 and 1.8. When generating multiple candidates, we utilize the vLLM framework (Kwon et al., 2023) that enables fast inference. To answer research questions related to candidate generation and reranking, we present the results from CodeLlama-7B-Instruct.
3.3 Reranking
For $n$-best reranking, we analyze three metrics types: likelihood-based, execution-based, and neural. For likelihood-based metrics, we use the log-likelihood of the generated generations given the input $x$ as well as CoderReviewer (Zhang et al., 2022) using $\log(p(x|y))+\log(p(y|x))$. [^4] In practice, we use the likelihood normalized by sequence length, as it mitigates bias towards shorter sequences. For execution-based metrics, we consider passing the trial unit test case as mentioned in Equation 2. For neural metrics, we use CodeScore (Dong et al., 2023) as it is so far the only publicly available neural metric that does not rely on execution during inference. [^5] We only test CodeScore on HumanEval since it was also trained on the original MBPP dataset. We do not consider LEVER (Ni et al., 2023) as it requires execution during inference time and requires the reranker to be trained on the target task.
For MBR, we consider execution-based metrics and neural metrics. Specifically, among execution-free metrics, we include CodeBertScore (Zhou et al., 2023), which is not trained with execution-based signals, and CodeScore (Dong et al., 2023), which is trained with execution-based signals. For execution-based metrics, we consider MBR-Exec (Shi et al., 2022) mentioned in Equation 4. Note that we assume access to the inputs of unit test cases used for evaluation, which aligns with one setting of MBR-Exec (Shi et al., 2022) and the setting of LEVER (Ni et al., 2023) and Self-Debug (Chen et al., 2024). We further study the effect of the number of test cases using EvalPlus (Liu et al., 2023), which contains over 100 high-quality unit test cases. Our setting is different from the setting using generated unit test cases (Chen et al., 2023; To et al., 2024; Huang et al., 2024) where normally neither trial unit test cases nor inputs of test cases for evaluation are available for either prompting or reranking. We leave the unification of these two settings to future work.
3.4 Self-Debugging
We consider the simple setting of self-debugging proposed by Chen et al. (2024). We consider the setting of using only unit test (UT) feedback, i.e., the feedback obtained from the compiler when a generated candidate is tested on trial unit tests. We only consider this simple setting because 1) our main focus is to test generally the impact of self-debugging on improving the oracle’s performance, 2) it does not require steps of generation using the LLM other than self-debugging, thus requires less computation for inference and 3) it is the feedback that gives the largest gain to execution accuracy post-debugging according to Chen et al. (2024). We perform 3 rounds of self-debugging for candidates generated with sampling because Chen et al. (2024) demonstrated that most debugging can be finished in 3 rounds. In each round, the inputs to the LLM include few-shot examples, the refined candidate from the previous round, and the candidate’s execution feedback on trial unit tests.
(a) HumanEval+
(b) MBPP-S+
Figure 3: Performance of MBR-Exec with and without filtering over different numbers of generated candidates using CodeLlama-7B-Instruct. We also provide the oracle with solid lines, denoted as Pass@k. Note that results that end with $+$ mean that it is evaluated on the plus with extended test cases. Results are averaged across at least 4 runs.
4 Results
In this section, we provide results and discuss the effects of reranking and self-debugging. Firstly, we analyze the impact of candidate generation on the model’s oracle performance (assuming we pick the best possible candidate) and the performance of execution-based reranking methods. We then compare different reranking methods and discuss the importance of execution for reranking. With the answers to these questions in mind, we study the effect of self-debugging and discuss the best way to combine all these methods to improve reranking performance. We provide a subset of all our experimental results to answer these research questions. For results on more experiments, please refer to Appendix A.
4.1 Candidate Generation
First, we study the impact of candidate generation on reranking. To do so, we vary the number of candidates sampled from CodeLlama-7B-Instruct, also exploring different temperature parameters. We also provide estimates of the oracle of reranking using Pass@k (Chen et al., 2021; Shi et al., 2022), commonly used as an evaluation metric for code generation models. Pass@k measures the proportion that at least one candidate passes all evaluation unit tests given $k$ generated candidates. We evaluate the performance of reranking on EvalPlus (Liu et al., 2023) for rigorous evaluation.
(a) HumanEval+
(b) MBPP-S+
Figure 4: Performance of MBR-Exec with and without filtering over sampling temperatures using CodeLlama-7B-Instruct with 50 generated candidates. We also provide the oracle of reranking with dashed lines. Note that results that end with $+$ mean that it is evaluated on the plus with extended test cases. Results are averaged across 4 runs.
What is the impact of the number of candidates?
We generate candidates using CodeLlama-7B-Instruct by setting temperature 1.6 and p 0.95. We hypothesize that filtering based on trial unit tests gives reranking better alignment to the ground-truth function regarding execution results.
Figure 3 shows the performance as a function of the number of generated candidates for different reranking approaches (based on execution) and the oracle performance for both HumanEval and MBPP-S. First of all, when filtering is not included, the performance of MBR-Exec only shows moderate improvement with more generated candidates across different numbers of unit test cases for evaluation and two benchmarks, which aligns with the analysis of Shi et al. (2022). However, with filtering, MBR-Exec shows considerable improvements over the version without filtering, and performance scales much better as more candidates are considered, which aligns with our hypothesis that filtering based on trial unit tests gives reranking better alignment to the ground-truth function regarding execution results.
Moreover, the performance of MBR-Exec with filtering consistently approaches the oracle performance even as we increase the number of candidates, which gives further evidence to our hypothesis that filtering based on trial unit tests gives reranking better alignment to the ground-truth function regarding execution results. Additionally, MBR with filtering has closer-to-oracle performance on the benchmark of MBPP-S tested on different setups with unit tests (see Figure 19(d)), while MBR with filtering approaches the oracle performance on both datasets (see Figure 19(c)), the gap is larger on the HumanEval dataset. This suggests that the overlaps of testing ability between trial unit tests and unit tests used for final evaluation are positively correlated. In contrast, 80% of the trial test cases are either included in the original test cases or the extended test cases from EvalPlus (Liu et al., 2023).
What is the impact of sampling temperature?
We compare generations with temperatures from 0.2 to 2.0 using CodeLlama-7B-Instruct with $p$ 0.95. We show results with 50 generated candidates in Figure 4. Note that previous studies on the optimal temperature are either based on worse-performing language models(Li et al., 2022), a different partition of the dataset (Shi et al., 2022), or a setting with trial unit tests not provided with the input (Chen et al., 2023).
When studying the effect of sampling temperature on the open-sourced CodeLlama-7B-Instruct and including the trial unit test in the prompt, we find that the temperature that gives the higher oracle for reranking is much higher than previously seen, with performance peaking at around 1.6 (while previous studied finding values <1 (Chen et al., 2021; Li et al., 2022; Shi et al., 2022; Chen et al., 2023)). We hypothesize that this results from better models allowing the generation of diverse and high-quality code.
Additionally, we observe a substantial difference in reranking performance with increased temperature when filtering based on the trial unit test included, for which we highlight the importance of filtering when reranking is considered. Finally, we adopt sampling temperature 1.6 with nucleus $p$ 0.95 for later experiments. For DeepSeekCoder-6.7B-Instruct, we use temperature 1.2.666With temperature ¿1.2 on DeepSeekCoder-6.7B-Instruct, sampling is more likely to generate token indices that are not defined in the vocabulary, see htatps://github.com/vllm-project/vllm/pull/3685.
HE (+) MBPP-S (+)
Random 34.1 (30.2) 50.1 (41.9)
Greedy 43.3 (39.0) 53.2 (44.8)
Oracle 82.4 (76.2) 83.2 (72.0)
N-Best Reranking
LL 45.1 (38.7) 54.2 (42.6)
CR 46.0 (40.5) 53.1 (43.1)
CS 33.1 (30.2) -
Filter 70.0 (59.7) 77.4 (60.8)
Filter + LL 71.0 (60.4) 78.1 (59.7)
Filter + CR 71.3 (61.3) 78.5 (60.2)
Filter + CS 71.2 (61.0) -
MBR
MBR-CBS 40.5 (35.5) 55.3 (45.4)
MBR-CS 37.5 (31.9) -
MBR-Exec 52.1 (49.5) 64.0 (57.8)
N-Best Reranking + MBR
LL + MBR-Exec 51.8 (49.4) 64.4 (58.1)
CR + MBR-Exec 51.4 (48.9) 64.8 (58.7)
Filter + MBR-CBS 69.8 (60.1) 77.3 (61.3)
Filter + MBR-CS 71.5 (59.9) -
Filter + MBR-Exec 74.5 (70.1) 79.0 (69.3)
Filter + LL + MBR-Exec 74.5 (70.6) 78.9 (69.2)
Filter + CR + MBR-Exec 74.0 (70.0) 79.3 (69.4)
Table 1: Performance across different reranking methods without and with filtering by sampling 50 candidates with temperature 1.6 and $p$ 0.95 using Code-Llama-7B-Instruct. For N-Best Reranking, we compare log-likelihood (LL), Coder-Reviewer (CR) (Zhang et al., 2022), and CodeScore (CS) Dong et al. (2023). For MBR, we compare CodeScore, CodeBertScore (CBS) (Zhou et al., 2023), and MBR-Exec (Shi et al., 2022). Note that HE stands for the HumanEval benchmark, $+$ refers to evaluation on extended test cases on EvalPlus. We also highlighted best and second best reranking results of the class of reranking methods.
4.2 Comparing Reranking Methods
With the best setting we observe from Section 4.1, we now try to provide an apples-to-apples comparison between different reranking methods that either directly utilize execution, predict executions, or utilize other information to understand the effect of filtering or execution more generally in reranking and provide suggestions for reranking methods.
We conduct experiments with $n$-best reranking and MBR to compare different raking approaches with and without filtering based on trial unit tests. We show our results in Table 1.
Filtering is the most powerful method for $n$-best reranking.
According to the group “N-Best Reranking” (see Table 1), filtering based on trial unit tests gives the largest gain from the random baseline compared to MAP decoding, CoderReviewer, and CodeScore. Note that while CodeScore is trained to predict the execution accuracy of generated code, it does not even outperform the random baseline, suggesting the gap in training execution-free rerankers for execution-based metrics.
Not only is filtering based on trial unit tests powerful in itself, but it also boosts the performance of other $n$-best reranking methods substantially. Specifically, we find that all methods including MAP decoding, CoderReviewer, and CodeScore outperform the filtering baseline moderately when combined with filtering except for the MAP decoding on MBPP-S+. Given our findings, we again emphasize the importance of filtering based on unit tests for reranking, which is a statement that is missing from previous studies that involve reranking (Shi et al., 2022; Zhang et al., 2022; Ni et al., 2023; Chen et al., 2024). Additionally, we suggest training better execution-free metrics to approximate reranking performance with filtering based on trial unit tests.
MBR-Exec with filtering still rocks.
According to Zhang et al. (2022), Coder-Reviewer outperforms MBR-Exec when filtering means if the candidate can give outputs based on the trial unit test inputs, and when MBR-Exec is only implemented on the trial unit tests. However, in our context based on a more sensible definition of filtering and MBR-Exec utilizing all unit test inputs, we found that both MAP and CoderReviewer underperform MBR-Exec with filtering. Additionally, when combined with filtering and MBR-Exec, they do not necessarily outperform MBR-Exec with filtering either, suggesting that the information from log-likelihood and “Reviewer” according to Zhang et al. (2022) is not necessarily helping reranking when sufficient direct execution is applied. Note that the MBR-Exec with filtering only slightly underperforms the oracle, suggesting there is little space to propose methods to outperform existing metrics that rely on both filtering and MBR-Exec. Instead, we suggest training a metric that does not rely on execution during inference is a better direction since MBR-Exec is also expensive with our current experimental setting.  [^7] We find difficulties in implementing multi-thread execution to get execution outputs, especially when numerous unit test cases are included (Liu et al., 2023). E.g. CodeLlama-7B-Instruct on HumanEval, generation takes around 3 hours while getting execution outputs takes around 5 hours.
4.3 Efficient Reranking with Fewer Unit Tests
While MBR-Exec performs well with filtering, it requires access to execution unit test inputs, which is impractical considering the execution cost. Therefore, we ask: how many unit tests should we execute candidates on? With this question in mind, we experiment with 50 generated candidates using MBR-Exec using increasing unit tests. Note that as we use the extended unit tests (Liu et al., 2023), we should also evaluate cases with the extended unit tests. Additionally, when the number of unit tests is lower than those given in the original dataset, we only use unit tests from the original dataset instead of the extended partition. We present results in Figure 5.
(a) HumanEval+
(b) MBPP-S+
Figure 5: Performance of MBR-Exec with fewer unit tests. We only test on HumanEval+ and MBPP+. Note that the original HumanEval benchmark contains an average of 9.6 unit tests per task, and the original MBPP benchmark contains around 3 unit tests per task.
With only a few more unit tests than from the original partition, MBR-Exec with filtering reaches peak performance.
The original HumanEval benchmark contains an average of 9.6 unit tests per task, and the original MBPP benchmark contains around 3 unit tests per task. Therefore, for HumanEval+, we only need 20 unit test cases, while for MBPP-S+, MBR-Exec with 5 unit test cases already gives optimal performance. This suggests that MBR-Exec can already reach its optimal performance with a few unit tests provided. Moreover, we again see the power of bringing in filtering based on trial unit tests, as it substantially improves reranking results across different benchmarks. Specifically, filtering brings a larger impact on HumanEval+ than MBPP-S+, where there’s a larger number of unit tests per task and a lower overlap between trial unit tests and unit tests used for evaluation, suggesting that they potentially have higher testing coverage compared to the trial unit tests provided in MBPP-S+.
(a) HumanEval+
(b) MBPP-S+
Figure 6: Improvement in Pass@k of CodeLlama-7B-Instruct after self-debugging compared to no self-debugging applied.
4.4 Improving the Oracle with Self-Debugging
Model HumanEval(+) MBPP-S(+)
MBR-Exec SD-1 SD-Multi MBR-Exec SD-1 SD-Multi
$\sim$7B Scale
CL-7B-Instruct 52.1 (49.5) 52.0 (49.5) 52.4 (50.0) 64.0 (57.8) 65.8 (59.1) 66.3 (60.4)
   + Filtering 74.5 (70.1) 74.4 (70.1) 76.5 (72.3) 79.0 (69.3) 79.3 (69.2) 81.3 (72.1)
DS-6.7B-Instruct 86.4 (85.5) 87.0 (86.0) 87.7 (87.0) 84.0 (78.6) 84.6 (78.8) 84.9 (79.3)
   + Filtering 91.6 (90.7) 91.5 (90.4) 91.6 (90.4) 89.1 (82.1) 89.0 (81.7) 89.9 (82.8)
$\sim$13B Scale
CL-13B-Instruct 61.3 (59.6) 61.7 (60.1) 62.7 (61.4) 73.9 (67.2) 74.4 (67.0) 74.4 (68.4)
   + Filtering 80.5 (76.1) 80.9 (76.5) 81.6 (77.6) 84.1 (74.7) 83.9 (74.2) 84.9 (76.2)
Table 2: Comparison of self-debugging methods with 50 candidates generated by CodeLlama-{7,13}B-Instruct with temperature 1.6 and DeepSeekCoder-6.7B-Instruct with temperature 1.2, and debugged over one (Chen et al., 2024) or multiple candidates using the same LLM. We also provide the upper bound after debugging. Note that results ending with +++ mean it’s evaluated on the plus with extended test cases. Results are averaged across 4 runs.
Since applying MBR-Exec with filtering based on trial unit tests approaches the theoretical oracle estimated by Pass@k, increasing this oracle seems more sensible than proposing better reranking methods. In this section, we aim to understand the effect of self-debugging on improving the oracle of performance and the performance of reranking generally. We then compare our proposed method of performing SD-Multi to just using one candidate. Note that for SD-Multi, we only use results with one round of debugging, while we use results with 3 rounds of debugging for SD-1.
Results suggest that self-debugging improves oracle performance of reranking, and one single round is enough.
According to Figure 6, we found that self-debugging does help to improve the oracle of reranking, represented by the improvement on Pass@k across different ks. However, later rounds of self-debugging do not show substantial improvement in the oracle compared to the first round.
SD-Multi outperforms SD-1.
According to Table 2, SD-Multi consistently outperforms SD-1 in most experiments, with better-performing models having smaller margins of improvement with self-debugging. The only exception we find is on experiments with candidates generated and self-debugged with DeepSeekCoder-6.7B-Instruct, where we find that the performance of reranking without self-debugging is already over 91, for which we presume that there is little space for improvement.
(a) HumanEval+
(b) MBPP-S+
Figure 7: Comparison of self-debugging methods over different numbers of candidates generated by CodeLlama-7B-Instruct, and debugged over $\{1,\textbf{Multi}\}$ candidates using the same LLM. We also provide the oracle after self-debugging for one round. Note that results ending with $+$ mean it is evaluated on the plus with extended test cases. Results are averaged across at least 4 runs.
Filtering is important in the context of self-debugging.
With 50 candidates generated and self-debugged, more experiments of MBR-Exec outperforming SD-1 with filtering compared to without filtering (see Table 2), suggesting that performance improvement brought by self-debugging can be partly accomplished by filtering, a finding that is missing from Chen et al. (2024) where filtering is also applied but not analyzed. Additionally, when filtering is applied, SD-1 only outperforms MBR-Exec with samples generated and debugged with CodeLlama-13B-Instruct.
The results of Figure 7 show that our proposed method consistently outperforms SD-1 with more rounds of debugging and that with 25 candidates generated, one-time self-debugging is more effective than sampling 50 candidates and debugging on a single candidate obtained by MBR-Exec. The results of MBPP are more moderate than the results of HumanEval, and we acknowledge that filtering closes the gap to the oracle, giving little space for improvement. Last but not least, self-debugging leaves a larger gap between our methods, with state-of-the-art results, and the oracle.
5 Related Work
Large Language Models for Programming Languages.
Recent years have witnessed the success of pretrained large language models for code generation (Chen et al., 2021; Li et al., 2022, 2023) that have been successful on multiple execution-based benchmarks (Hendrycks et al., 2021; Austin et al., 2021; Chen et al., 2021; Liu et al., 2023; Lai et al., 2022; Wang et al., 2022). More recently, strong-performing open-source pretrained Code LLMs (Rozière et al., 2024; Luo et al., 2024; Guo et al., 2024a; Lozhkov et al., 2024) have been available, and they can perform tasks such as filling-in-the-middle (Rozière et al., 2024), code debugging (Guo et al., 2024b), as some are also trained with instruction tuning. Our works use open-source code LLMs and instruction tuning to study reranking and self-debugging.
Reranking and MBR Decoding.
Reranking and MBR Decoding have been studied extensively in domains such as machine translation (Eikema and Aziz, 2020, 2022; Fernandes et al., 2022; Farinhas et al., 2023). Similarly, Shi et al. (2022) and Li et al. (2022) proposes MBR decoding based on agreement on execution results, with Li et al. (2022) further propose filtering out samples that do not pass the trial unit tests. Shi et al. (2022) propose filtering out samples that cannot get execution outputs, followed by Zhang et al. (2022) who combine this filtering with the log-likelihoods of the generated candidate, and the log-likelihoods of input given the generated candidate. Our work differs from these works as we report the effect of filtering based on trial unit tests on candidate generation, reranking, and self-debugging. Our work also differs in combining reranking with self-debugging.
Other works proposed methods with generated unit test cases with Chen et al. (2023) and To et al. (2024) utilizing generated unit tests for reranking and Huang et al. (2024) incorporating both generated unit test cases and execution-based specifications. Our work differs from this line of research as we adopt the setting with access to trial unit tests and inputs of all unit tests for evaluation to study both filtering based on trial unit tests and MBR-Exec on multiple test cases, while Chen et al. (2023) and To et al. (2024) presume access to neither for prompting or reranking and Huang et al. (2024) only incorporate trial unit tests in some experiments without directly applying filtering based on them. While Chen et al. (2023) experiment with filtering in other algorithmic coding tasks in the appendix, our work differs in showing the effect of filtering on candidate generation, efficient reranking with fewer test cases, and self-debugging. Additionally, our work differs in that we experiment with different candidate generation settings with filtering being applied that leads to findings of new candidate generation “sweet spots”. Our work also differs in combining reranking with self-debugging.
Additionally, previous works have also studied training N-Best rerankers. Early efforts include training classifiers that predict program success in execution (Inala et al., 2022), followed by Ni et al. (2023) proposing to train both model-specific and task-specific classifiers for reranking. Our work differs from Ni et al. (2023) as we evaluate with more general-purpose metrics (Dong et al., 2023) that do not rely on execution during inference time. Moreover, we include filtering based on trial unit tests as part of the feedback, our focus is on comparing execution-based and execution-less metrics. The redundancy of using trained rerankers on target tasks and target code generation models with execution included even during inference time is also supported by our results of filtering with MBR-Exec.
Last but not least, our work differs from all these works as we also study the neural metrics (Zhou et al., 2023; Dong et al., 2023) in comparison to execution-based metrics.
Self-Debugging with LLMs
Chen et al. (2024) study the effect of using the same Code LLM to debug a single generated code candidate by varying feedback information such as execution errors and code explanation. Olausson et al. (2024) also study self-debugging but focus on different sampling methods and their impacts on self-debugging. Zhong et al. (2024) additionally integrate information regarding intermediate states to provide additional information for debugging. Our work differs from theirs as we study the joint effect of both self-debugging and reranking. We also perform self-debugging on multiple candidates.
6 Conclusion and Future Work
We leverage recent advances in code generation, including N-best reranking, MBR, and self-debugging and we propose DOCE, a unified framework for generating code with higher execution accuracy. We perform an extensive study with our proposed framework on reranking by studying candidate generation, comparing across reranking methods, and studying self-debugging for increasing reranking upper bound. Importantly, we highlight the impact of filtering based on trial unit tests, a commonly used technique whose effect has been overlooked in previous works, showing that it is crucial for all procedures DOCE. Last but not least, we observe that self-debugging improves the upper bound, with self-debugging on all generated candidates giving further improvement of the reranking results compared to self-debugging on one single candidate. Potential future works include incorporating studying the effects of generated unit tests and specifications, studying the effect of directly and indirectly using feedback from trial unit tests, and training better general-purpose execution-less metrics for reranking.
Limitation
First of all, our paper is limited by the scale of experiments on more code generation models, due to the limit of compute and the limit. Our solution for this is to select representative classes of open-source models and experiments on models with parameters below 15 billions.
Appendix A Full Results
A.1 Candidate Generation
We present the effect of the number of candidates with candidates generated by CodeLlama-{7,13}B-Instruct (see Figure 8 and Figure 10) DeepSeekCoder-6.7B-Instruct (see Figure 9) on both the original unit tests and extended test cases in EvalPlus (Liu et al., 2023). Most results align with those we present in Figure 3, where filtering based on trial unit tests improves MBR-Exec considerably and approaches the oracle performance when combined with MBR-Exec. The only exception happens on candidates generated by DeepSeekCoder-6.7B-Instruct on HumanEval (see Figure 9(a)) where the execution accuracy using MBR-Exec with filtering on no less than 10 candidates is already above 90.
To validate our choice of temperature, we also present the choice of sampling temperature of CodeLlama-7B-Instruct on both the original and extended test cases in Figure 11. For CodeLlama-13B-Instruct and DeepSeekCoder-6.7B-Instruct, we compare results with our choice of temperature for further experiments with results using temperature 0.8. Results are presented in Table 3. All models we experiment with allow a sampling temperature over 1, with lower mean execution accuracy but higher oracle performance. Combined with filtering on trial unit tests, MBR-Exec allows constant improvement in execution accuracy when sampling with higher temperatures, which is not guaranteed without filtering.
A.2 Comparing Reranking Methods
We compare reranking methods with candidates generated by DeepSeekCoder-6.7B-Instruct (see Table 4) and CodeLlama-13B-Instruct (see Table 5). Findings are similar to Section 4.2, where filtering based on trial unit tests is important itself and helps to boost other reranking methods, and MBR-Exec remains the best-performing method, giving close-to-oracle performance when combined with filtering. Log-likelihoods of the generated candidate and the input do not necessarily help better reranking.
A.3 MBR-Exec with Fewer Unit Tests
We present results using fewer unit tests using candidates generated from DeepSeekCoder-6.7B-Instruct Figure 12 and CodeLlama-13B-Instruct in Figure 13. Similar to Figure 5, we only need 20 unit test cases for HumanEval+, and 5 unit test cases for MBPP-S+ to obtain optimal performance with MBR-Exec.
A.4 Improving Oracle with Self-Debugging
We first present results of improvement of Pass@k that estimates oracle improvement using candidates generated and self-debugged by CodeLlama-7B-Instruct (see Figure 14), DeepSeekCoder-6.7B-Instruct (see Table 15) and CodeLlama-13B-Instruct (see Table 16). Our findings align with Section 4.4 as one round of self-debugging is enough to improve the oracle.
We then present results over different numbers of candidates generated and debugged by CodeLlama-7B-Instruct (see Figure 17), DeepSeekCoder-6.7B-Instruct (see Table 18) and CodeLlama-13B-Instruct (see Table 19). We find that SD-Multi outperforms SD-1 constantly except for HumanEval(+) with candidates generated and debugged by DeepSeekCoder-6.7B-Instruct with 25 generated candidates.
Appendix B Prompts
B.1 Prompt for Code Generation (0-Shot)
1[INST] Write code to solve the following coding problem that obeys the constraints and passes the example test cases. Please wrap your code answer using ‘‘‘:
2‘‘‘python
3{prompt}
4‘‘‘
5[/INST]
6‘‘‘python
Figure 20: 0-shot prompt for code generation with CodeLlama-{7,13}B-Instruct.
1You are an AI programming assistant, utilizing the DeepSeek Coder model, developed by DeepSeek Company, and you only answer questions related to computer science. For politically sensitive questions, security and privacy issues, and other non-computer science questions, you will refuse to answer.
2### Instruction:
3Please complete the following Python function in a markdown-style code block:
4‘‘‘python
5{prompt}
6‘‘‘
7### Response:
8‘‘‘python
Figure 21: 0-shot prompt for code generation with DeepSeekCoder-6.7B-Instruct.
See Figure 20 and Figure 21.
B.2 Prompt for Reviewer (3-Shot)
1### task start ###
2### code
3‘‘‘python
4def count_ways(n):
5 A = [0] * (n + 1)
6 B = [0] * (n + 1)
7 A[0] = 1
8 A[1] = 0
9 B[0] = 0
10 B[1] = 1
11 for i in range(2, n+1):
12 A[i] = A[i - 2] + 2 * B[i - 1]
13 B[i] = A[i - 1] + B[i - 2]
14 return A[n]
15‘‘‘
16
17Write the docstring for the above code.
18
19### docstring
20‘‘‘python
21"""
22Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
23$>$ count_ways(2)
243
25"""
26‘‘‘
27### task end ###
28
29### task start ###
30### code
31‘‘‘python
32import re
33def find_char_long(text):
[…]
35‘‘‘
36
37Write the docstring for the above code.
38
39### docstring
40‘‘‘python
41"""
42Write a function to find all words which are at least 4 characters long in a string by using regex.
43$>$ find_char_long(’Please move back to stream’)
44[’Please’, ’move’, ’back’, ’stream’]
45"""
46‘‘‘
47### task end ###
48
49### task start ###
50### code
51‘‘‘python
52def square_nums(nums):
53 square_nums = list(map(lambda x: x ** 2, nums))
54 return square_nums
55‘‘‘
56
57Write the docstring for the above code.
58
59### docstring
60‘‘‘python
61"""
62Write a function to find squares of individual elements in a list using lambda function.
63$>$ square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])
64[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
65"""
66‘‘‘
67### task end ###
1### task start ###
2### code
3‘‘‘python
4{code}
5‘‘‘
6
7Write the docstring for the above code.
8
9### docstring
10‘‘‘python
11{prompt}
12‘‘‘
Figure 22: 3-shot prompt for getting Reviewer log-likelihoods (Zhang et al., 2022) on HumanEval.
1### task start ###
2### code
3‘‘‘python
4def count_ways(n):
5 A = [0] * (n + 1)
6 B = [0] * (n + 1)
7 A[0] = 1
8 A[1] = 0
9 B[0] = 0
10 B[1] = 1
11 for i in range(2, n+1):
12 A[i] = A[i - 2] + 2 * B[i - 1]
13 B[i] = A[i - 1] + B[i - 2]
14 return A[n]
15‘‘‘
16
17Write the docstring for the above code.
18
19### docstring
20‘‘‘python
21"""
22Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
23assert count_ways(2) == 3
24"""
25‘‘‘
26### task end ###
27
28### task start ###
29### code
30‘‘‘python
31import re
32def find_char_long(text):
[…]
34‘‘‘
35
36Write the docstring for the above code.
37
38### docstring
39‘‘‘python
40"""
41Write a function to find all words which are at least 4 characters long in a string by using regex.
42assert find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]
43"""
44‘‘‘
45### task end ###
1### task start ###
2### code
3‘‘‘python
4def square_nums(nums):
5 square_nums = list(map(lambda x: x ** 2, nums))
6 return square_nums
7‘‘‘
8
9Write the docstring for the above code.
10
11### docstring
12‘‘‘python
13"""
14Write a function to find squares of individual elements in a list using lambda function.
15assert square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])==[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
16"""
17‘‘‘
18### task end ###
19
20### task start ###
21### code
22‘‘‘python
23{code}
24‘‘‘
25
26Write the docstring for the above code.
27
28### docstring
29‘‘‘python
30{prompt}
31‘‘‘
Figure 23: 3-shot prompt for getting Reviewer log-likelihoods (Zhang et al., 2022) on MBPP.
See Figure 22 and Figure 23.
B.3 Prompt for Self-Debugging (6-Shot)
1### task start ###
2### prompt
3‘‘‘python
4def count_ways(n):
5 """ Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
6 $>$ count_ways(2)
7 3
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def count_ways(n):
14 if n == 0:
15 return 1
16 if n == 1:
17 return 1
18 if n == 2:
19 return 3
20 return count_ways(n-1) + count_ways(n-2)
21‘‘‘
22
23### feedback
24With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
25The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
26
27### fixed code
28‘‘‘python
29def count_ways(n):
30 A = [0] * (n + 1)
31 B = [0] * (n + 1)
32 A[0] = 1
33 A[1] = 0
34 B[0] = 0
35 B[1] = 1
36 for i in range(2, n+1):
37 A[i] = A[i - 2] + 2 * B[i - 1]
38 B[i] = A[i - 1] + B[i - 2]
39 return A[n]
40‘‘‘
41
42### feedback
43With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
44The code above is correct.
45### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def differ_At_One_Bit_Pos(lhs,rhs):
5 """ Write a python function to check whether the two numbers differ at one bit position only or not.
6 $>$ differ_At_One_Bit_Pos(15,8)
7 False
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def differ_At_One_Bit_Pos(lhs,rhs):
14 """ Write a python function to check whether the two numbers differ at one bit position only or not.
15 $>$ differ_At_One_Bit_Pos(15,8)
16 False
17 """
18 if (lhs - rhs) == 0 or (lhs - rhs) == 1:
19 return True
20 return False
21‘‘‘
22
23### feedback
24With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
25The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
26
27### fixed code
28‘‘‘python
29def differ_At_One_Bit_Pos(a,b):
30 """ Write a python function to check whether the two numbers differ at one bit position only or not.
31 $>$ differ_At_One_Bit_Pos(15,8)
32 False
33 """
34 x = a ^ b
35 return x and (not(x & (x - 1)))
36‘‘‘
37
38### feedback
39With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
40The code above is correct.
41### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def find_char_long(text):
5 """ Write a function to find all words which are at least 4 characters long in a string by using regex.
6 $>$ find_char_long(’Please move back to stream’)
7 [’Please’, ’move’, ’back’, ’stream’]
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def find_char_long(text):
[…]
15‘‘‘
16
17### feedback
18With the above function, theres following error:
19"""
20NameError: name ’re’ is not defined
21"""
22The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
23
24### fixed code
25‘‘‘python
26import re
27def find_char_long(text):
[…]
29‘‘‘
30
31### feedback
32With the above function, the assertion is find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]‘ and the real execution output is ‘[’Please’, ’move’, ’back’, ’stream’]‘. The code passes the assertion.
33The code above is correct.
34### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def square_nums(nums):
5 """ Write a function to find squares of individual elements in a list using lambda function.
6 $>$ square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])
7 [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def square_nums(nums):
14 """ Write a function to find squares of individual elements in a list using lambda function.
15 $>$ square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])
16 [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
17 """
18 square_nums = list(map(lambda x: x ** 2, nums))
19 return square_nums
20‘‘‘
21
22### feedback
23With the above function, the assertion is square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10]) == [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘ and the real execution output is ‘[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘. The code passes the assertion.
24The code above is correct.
25### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def find_Rotations(s):
5 """ Write a python function to find the minimum number of rotations required to get the same string.
6 $>$ find_Rotations("aaaa")
7 1
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def find_Rotations(s):
14 if len(s) == 1:
15 return 1
16 elif len(s) == 0:
17 return 0
18 else:
19 l = len(s)
20 min = len(s)
21 for i, c in enumerate(s):
22 if c == s[0]:
23 temp = i
24 if temp < min:
25 min = temp
26 return min
27‘‘‘
28
29### feedback
30With the above function, the assertion is find_Rotations("aaaa") == 1‘ but the real execution output is ‘0‘.
31The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
32
33### fixed code
34‘‘‘python
35def find_Rotations(s):
36 n = len(s)
37 for i in range(1, n + 1):
38 if s[i:n] + s[0:i] == s:
39 return i
40 return n
41‘‘‘
42
43### feedback
44With the above function, the assertion is find_Rotations("aaaa") == 1‘ and the real execution output is ‘1‘. The code passes the assertion.
45The code above is correct.
46### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def small_nnum(lst,n):
5 """ Write a function to get the n smallest items from a dataset.
6 $>$ small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2)
7 [10, 20]
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def small_nnum(lst,n):
14 """ Write a function to get the n smallest items from a dataset.
15 $>$ small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2)
16 [10, 20]
17 """
18 lst = sorted(lst)
19 lst = lst[:n]
20 return lst
21‘‘‘
22
23### feedback
24With the above function, the assertion is small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2) == [10,20]‘ and the real execution output is ‘[10, 20]‘. The code passes the assertion.
25The code above is correct.
26### task end ###
1### task start ###
2### instruction
3{instruction}
4‘‘‘python
5{unit_test}
6‘‘‘
7
8### code
9‘‘‘python
10{code}
11‘‘‘
12
13###feedback
14{feedback}
15The code above is
Figure 24: 6-shot prompt for self-debugging on HumanEval.
1### task start ###
2### instruction
3Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
4‘‘‘python
5assert count_ways(2) == 3
6‘‘‘
7
8### code
9‘‘‘python
10def count_ways(n):
11 if n == 0:
12 return 1
13 if n == 1:
14 return 1
15 if n == 2:
16 return 3
17 return count_ways(n-1) + count_ways(n-2)
18‘‘‘
19
20### feedback
21With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
22The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
23
24### fixed code
25‘‘‘python
26def count_ways(n):
27 A = [0] * (n + 1)
28 B = [0] * (n + 1)
29 A[0] = 1
30 A[1] = 0
31 B[0] = 0
32 B[1] = 1
33 for i in range(2, n+1):
34 A[i] = A[i - 2] + 2 * B[i - 1]
35 B[i] = A[i - 1] + B[i - 2]
36 return A[n]
37‘‘‘
38
39### feedback
40With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
41The code above is correct.
42### task end ###
1### task start ###
2### instruction
3Write a python function to check whether the two numbers differ at one bit position only or not.
4‘‘‘python
5assert differ_At_One_Bit_Pos(15,8) == False
6‘‘‘
7
8### code
9‘‘‘python
10def differ_At_One_Bit_Pos(lhs,rhs):
11 if (lhs - rhs) == 0 or (lhs - rhs) == 1:
12 return True
13 return False
14‘‘‘
15
16### feedback
17With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
18The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
19
20### fixed code
21‘‘‘python
22def differ_At_One_Bit_Pos(a,b):
23 x = a ^ b
24 return x and (not(x & (x - 1)))
25‘‘‘
26
27### feedback
28With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
29The code above is correct.
30### task end ###
1### task start ###
2### instruction
3Write a function to find all words which are at least 4 characters long in a string by using regex.
4‘‘‘python
5assert find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]
6‘‘‘
7
8### code
9‘‘‘python
10def find_char_long(text):
[…]
12‘‘‘
13
14### feedback
15With the above function, theres following error:
16’’
17NameError: name re is not defined
18’’
19The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
20
21### fixed code
22‘‘‘python
23import re
24def find_char_long(text):
[…]
26‘‘‘
27
28### feedback
29With the above function, the assertion is find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]‘ and the real execution output is ‘[’Please’, ’move’, ’back’, ’stream’]‘. The code passes the assertion.
30The code above is correct.
31### task end ###
1### task start ###
2### instruction
3Write a function to find squares of individual elements in a list using lambda function.
4‘‘‘python
5assert square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])==[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
6‘‘‘
7
8### code
9‘‘‘python
10def square_nums(nums):
11 square_nums = list(map(lambda x: x ** 2, nums))
12 return square_nums
13‘‘‘
14
15### feedback
16With the above function, the assertion is square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10]) == [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘ and the real execution output is ‘[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘. The code passes the assertion.
17The code above is correct.
18### task end ###
1### task start ###
2### instruction
3Write a python function to find the minimum number of rotations required to get the same string.
4‘‘‘python
5assert find_Rotations("aaaa") == 1
6‘‘‘
7
8### code
9‘‘‘python
10def find_Rotations(s):
11 if len(s) == 1:
12 return 1
13 elif len(s) == 0:
14 return 0
15 else:
16 l = len(s)
17 min = len(s)
18 for i, c in enumerate(s):
19 if c == s[0]:
20 temp = i
21 if temp < min:
22 min = temp
23 return min
24‘‘‘
25
26### feedback
27With the above function, the assertion is find_Rotations("aaaa") == 1‘ but the real execution output is ‘0‘.
28The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
29
30### fixed code
31‘‘‘python
32def find_Rotations(s):
33 n = len(s)
34 for i in range(1, n + 1):
35 if s[i:n] + s[0:i] == s:
36 return i
37 return n
38‘‘‘
39
40### feedback
41With the above function, the assertion is find_Rotations("aaaa") == 1‘ and the real execution output is ‘1‘. The code passes the assertion.
42The code above is correct.
43### task end ###
1### task start ###
2### instruction
3Write a function to get the n smallest items from a dataset.
4‘‘‘python
5assert small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2) == [10,20]
6‘‘‘
7
8### code
9‘‘‘python
10def small_nnum(lst,n):
11 lst = sorted(lst)
12 lst = lst[:n]
13 return lst
14‘‘‘
15
16### feedback
17With the above function, the assertion is small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2) == [10,20]‘ and the real execution output is ‘[10, 20]‘. The code passes the assertion.
18The code above is correct.
19### task end ###
1### task start ###
2### instruction
3{instruction}
4‘‘‘python
5{unit_test}
6‘‘‘
7
8### code
9‘‘‘python
10{code}
11‘‘‘
12
13###feedback
14{feedback}
15The code above is
Figure 25: 6-shot prompt for self-debugging on MBPP-S.
See Figure 24 and Figure 25.